-- MarkWard - 08 Mar 2017


  • Installed 3 of the 4 - 4 shooter cameras
  • (Means all 4 cameras are now on DMTPC)
  • Leak checks of the vessel look good, should still be monitored.
  • Light leaks - Will will have more information but, looks ok, there was a small leak on the bottom right camera due to a poor gasket, fixed now using an Al blanket.
  • Started to mount the Cathode onto the frame - still have some alignment issues, need to check the drawings etc
  • Cathode being mounted should allow us to start fixing the TPC and inserting the resistance chain, still have issues with it falling into an egg shape.
  • Started looking at LED light source smile


  • Gas is still holding us up - tried to get some from the biological science dept but, the contact person was not there
  • Top right camera is installed, focussed to 70cm, might need to adjust it based on an e-mail from Gabriela
  • Extra light tightening has been done using Al tape and foil.
  • Camera is being used to search for light leaks - Will.... will have more information
  • DAQ PC has been moved downstairs - operational
  • Digitizers have been moved downstairs - to be tested
  • Vessel leak check for the top right is signed off chamber ultimate pressure is approximately 1.8e-8 mbar, it fluctuates with the room temperature


  • More fixes on m3slow, correct DAQ page now appears when pressing the link, it opens in its own window though rather than with the menus, ok for now... IR not a HTML/Java/Web programmer person soooooo yeah, slowly slowly
  • Found a few bugs with the web interface will try and figure them out, Example, SynOP gives the ambiant temperature, yesterday it was showing 200+ C in the slow control page, but normal 21 in the SynOp interface.
  • Installed Top right camera port
  • Stalled by a couple of items
    • Waiting on compressed gas, the installation emptied our supply, two bottles have been ordered
    • Temp-fix involves a bike pump... we can in theory open and close valves, gate valve is 50/50
  • Ultimate pressure of the chamber overnight was 2.7E-8 mbar, which is close to what it was previously with a blank.
    • First RoR test gives 3.8x10^-6 mbar l / s compared to 2.8x10^-6 mbar l / s before the installation
    • Most probable from water/atmospherics introduced with the new flange.
    • TR status changed to leak checked
  • Appears we have a faulty/blocked fume hood, rise of temperature in the lab resulting from being unable to exhaust the aircon was causing the chamber pressure to rise.
    • Will investigate - could be a clean room hat stuck in the exhaust/birds/faulty fan?
  • Started hassling SAM for safety documents


  • Moved m3slow down to the lab - long story short... its working http://m3slow.pp.rhul.ac.uk/index.html
  • Placed a net-switch in the lab to fan-out the network
  • Reassisned the SynOp netbooter to its pp network ip address
  • Moved caen/NIM crate down to the lab
  • Started (and completed) getting slow control aspects of DMTPC wired in and read out
    • Pressure gauges
    • anode + cathode HV supplies <--- This one was the cause of all the errors that where killing the slow control, found out that not all the caen powersupplies we have play nicely over USB (some commands are different and cause fatal errors)
  • changes made on m3slow
    • Fixed broken links to Synop - now hard coded rather than using the iptables
    • Fixed broken links to DB to dump a table of all fields
    • Added a 'online' plot for the vacuum pressure
  • Tests performed
    • Start and stop slow control
    • HV ramping, and changing parameters
    • making gif-plots using the interface
    • dumping db table and parameters
    • db export to text (using scripts)
  • Things to be tested
    • Pressure control (right now im protecting/paranoid about the turbo, so will run initial tests with everything off)
  • Continued with pumpout and leak rate measurements
    • See Zach's update
    • Spoilers - looks good
    • Have run the TMP overnight (blanked off with the gate) - flawless operation so far touch wood taps on head


  • Took one for the team in baseball, GO TEAM STAFF! Minor bruising to thumb sustained :p, did not hinder the consumption of BBQ food or beverages in the team building activities that followed the game
  • IMG_20170526_170335.jpg
  • Bottle rack installed in the lab
  • Moved a N2 bottle downstairs and pressurized the valves, now in process of checking for minor leaks in the manifold that feeds the valves, will switch this for a cheaper pressure source ASAP
  • provision on this setup to drive the valves, might need to be very careful if we want to use it to back fill the chamber, BUT if there no leaks in the gas lines (shouldn't be anything major) then it is just a case of locking valves at the pressure they require and altering a valve. I will work on a better solution for this once we have more final bottles in place.
  • Test cycled valves we need for the pump down
  • Water is plumbed in, new hose clamps are installed, Mass away Tuesday and Today, Andy will help me get the water system going regardless
  • Next step for valve automation is to get the netbooter working on the network, for simplicity we should adapt power plugs a minimal amount of times so propose to use the netbooter (other than valves) as a US power strip for a number of smaller items
  • Meeting with Sam to discuss size changes to frame and movement of lifting hardware such that the chamber can be wheeled through the corridor and not require another window juggle.


  • Away last week getting grilled in Carleton
  • Parts for the gas system to valves in T111 arrived and have been deployed, most is in place
  • Frame for TPC construction has been constructed (see Zacks update)
  • Water system is almost ready to go parts and fittings are mostly together, needs finally connections made (was working on the gas side yesterday)
  • Electrical for visual readout (with controller) of the gauges is online. Still requires daq linkup
  • NI input controller is now located in the rack in T111 pointed toward the vessel.
  • Electrical box for valve PSUs is underconstruction (no upset health and safety people), task put in by Zack


  • TPC frame will not fit through any of the doors or in the lift, needs to be redesigned
  • Spent a few days redesigning the frame in AutoCAD, parts to be cut, going with the design, parts are with the workshop
  • frame1.png
  • Key Features
    • Two separate parts - Box and stand
    • Box is sized based on the chamber supports that hold the TPC
    • Means the TPC can be built and transported to the lab via the elevator... JUST (1cm to spare)
    • Once in the lab, the Box section can be attached to its stand, and the height adjusted.
    • The open side of the box along with the ability to slide out the supports mean the TPC can be introduced as close as possible to the vessel
    • The stand is designed not to interfere with the suppoting frame of the vessel.
  • Ball is rolling on our water cooling loop - will need only minimal parts ontop of what we have from MIT, 'should' be up and going soon™
  • Window fan is installed (finally)
  • Just got the 3d drawings of the vessel back from sam!... need to ask "are the attachments flexable? can we detach it easily and mout it on the skid for the pumps... looks like it would be an issue for off axis like this... unless we simulate protons/pions vs valves.
  • Untitled.png


  • T111 Cleaning - Floor needs an almost daily mop, it appears to gather fluff, not sure if its from the outside or is coming under the door.
  • Particle Counter - Keeping a paper chart (for now) of particle count in the lab in general, and on North (window facing) and South inside the tent. Tent is maintaining Class 7 or better spec for the measurements iv made. Depends heavily upon location and proximity to objects in the tent and when the floor was last cleaned.
  • Recommend investing on some slippers for the cleantent, the blue booties last approximately 30 seconds.
  • Installation of Vacuum services on DMTPC - Basic services are installed, Manometer, vacuum gauges on vessel and turbo, Gas inlet valve, turbo valve, gate valve... Basically hardware is in place except for final flex-hose connection between the vessel/turbo and the roughing pump.
  • Found some evidence of deposits or corrosion on the fanblades of the turbo, centered around one area
  • Image-1.png
  • Looks like it is either corrosion or dried/welded onto the surface. Light rubbing with damp kimwipes didn't do much, Alcohol was also ineffective.

  • Gate Valve and Turbo pump are obviously very heavy, thanks to everyone we drafted in to help, they both went on fairly easily, we should get/make a couple of thin metal prods for future installation to hold/move the copper gasket while marrying up the two sealing surfaces.
  • Power connector for the Roughing pump was given to Hakim, should already be done, the phases need to be checked, apparently the building has a couple of the phases switched around (Joy-Joy!) but Hakim is already well aware and will check.
  • Items to do:
    • Final vacuum connection
    • Water cooling connections - need to check spec and fittings, the chiller loop will already have a mould inhibitor so we shouldn't need to get any Glycol.
    • Electrical - Gauges and valves - worrying that the valve control is basically a stack of powersupplies in a cardboard box... Id recommend we mount these in a box, otherwise people will have a minor meltdown
    • Compressed Air for driving valves



  • Continued to check concentration in the case of bottle failures. Recap from last week, with the addition of a 15min leak for the cylinder. You can see that using only the fume hood, we dip below the alarm setpoint
  • O2_Conc_HudOnly.png
  • This doesn't include a circulation fan mounted in the window, not sure of the spec, but 12 inch window fans can move air at a rate up to around 1200m3/hr
  • Examine original case with one bottle, and add to it the requirements of two bottles.
  • O2_Conc_HudAndFan.png
  • Result is, if we have a window fan producing greater than 850m3 it is safe to house two bottles in the lab if they are both to fail
  • Probability of two bottle failures at the same time is extremely rare.

  • As everyone knows, DMTPC is now in T111
  • space is at a premium, need a lab cleaning day,,, this afternoon? Tomorrow?
  • Spent a couple of afternoons cleaning and organising the lab.


  • T111
    • Looked at the Oxygen deficency hazard for T111 in the case of useing gas bottles in the room.
    • BCGA Code of Practice 44 guidelines describes a bottle failing in some manner and dumping the whole contents in an extremely short amount of time is an extremely unlikely event, thus when concidering ODH, you should not assume this is the absolute worse case scenario
    • Guidelines contain differencial equation solutions for a room with fresh air turn over with a gas source leaking.
    • It goes as follows
    • O2Conc.png
    • For T111 if we assume Volume is approx 85m3, we can ask, what is the room turn over required not to trigger any kind of Oxygen alarm should a bottle empty its contence within an hour.
    • Standard Cylinder typically contains 10m3 of gas
    • What is the recovery.
    • We find the following.
    • O2Conc_trend.png
    • Looks like we need to turn over the gas volume in the room every hour, which requires volume flow of 24l/s, looks like a large number HOWEVER
    • Looking at the fume hood, the face velocity is 0.55m/s and the minimum intake area is 0.06m2, resulting in a volumetric flow of approx 40 l/s or 1.7 turn over / hr
    • If we add the new ventillation fan which will go into the window, the situation gets even better.
    • So it looks like it isn't a difficult argument to make that bottle use in the lab will not pose a health and safety risk.

    • Produced a CAD drawing for T111 to help plan locations and free space.
    • T111.png: All numbers are in mm, though the room is basically built in inches, and the layout isn't exactly stright, So there will be some minor differences between reality and the drawing.


  • IMG_20170327_094216.jpg

  • IMG_0336.JPG

  • The transport crew had an audience
  • Other shipments that arrived that day was the new clean tent, which was sent in a huge truck, with a driver who had been given no instructions from Cleanrooms.
    • Despite that, we got it offloaded as quickly as we could, all equipment is now stored in the teaching lab.
  • Thanks to everyone who helped move all the boxes inside! Was a fantastic effort, especially those who came back from Canada and came straight to work to help.
  • All boxes have been checked off, nothing appears lost or unaccounted for
  • Need a replacement 750W PSU for one of the computers note to everyone, always ensure PSUs can take 240V when they come from the US >_<

  • Other things :
    • Gathering prices for Flow Control devices from Omega and MKS to add to our quote from Alicat, I will collate them into a spreadsheet. Omega's prices are on their webpage, MKS just sent me some numbers about 10 minutes ago.


  • Attended the 2nd Workshop on Neutrino Near Detectors based on gas TPCs
    • Most discussion was about changes to the ND280 detector, lots of different proposals by the different groups
    • Most popular seemed to be Removal of the P0D, re-use of the ECals Possible upgrade to the P0D ECal, re-use of the TPCs and fill the extra space with TPCs and TOF systems
    • -Image
* nd280.png:
  • My Inquiries in regard to recertifying our gas board with BOC and HEROSE has been met with 100% the opposite interpretation by Andy,
    • Told the users where informed work needed to be done... but this is about it as far as I understand
    • Leaves the labs floating not knowing what the best way forward is. So when I was tasked to look into it after a couple of phone calls it would appear the maintainence schemes are written on the side of extreme paranoia, but legally there is no set dates for replacement of items in a high pressure system
    • Speck and Burke appear to have simply given all the systems the minimum lifespan, cynically, so they can sell the department lots of expensive equipment every 5 years. They are a provider do not actually do re-certification.
    • HEROSE say they can reset and certify at a significant cost reduction.
    • BOC representative said the equipment we have is 'high end' and in general for the regulators we have they'd put 5 years warrenty on them, which means we definitely don't need to replace them, and only have to perform an action because of the written scheme.
    • Basically it is my personal oppinion that the written scheme is on the side of extreme paranoia, and doesn't reflect the reality of our system and its safe operation. - this is my oppinion based on - experience and speaking with BOC and HEROSE.
    • Default plan for the department was "Buy all new equipment" and no one questioned the 5 year lifetime of very expensive equipment.
    • Quite frustrating, we appear at the mercy of what ever supplier we go with, which looks to simply be expensive.
    • Quite frustrating... Interesting to be told that A company willing to certify things, wouldn't meet legal requirements or be insurable... Id rather not say much more.
    • Frustrating. hits head on table
  • Delivery of our MIT shipment should be happening on Friday, unless Ian finds out to the contrary.
    • Initial plan would be to deliver all 7 boxes on Friday, the time could be anytime within working hours... we might need a team of people to help get the smaller boxes unpacked and into the lab.
    • The vessel wont be lifted into the building on Friday, this is almost a certenty, plan would be to ensure it is weather sealed (Packing of the vessel itself looks very good for weather sealing, but we would do some more just to be sure.)


  • University ordered a shutdown of unserviced gas systems, untill services/repalcements can be made (Safety!)
  • GasCabinet.png:
  • Searching for solution to regulator servicing/replacement (Gas cabinet in Dark Matter Lab)
    • Only advancement is that HEROSE will reset and service the main relief valves in the system
    • Approx 20 per valve for service vs 100-150 per valve
    • BOC are yet to quote on service of regulartors frown
  • Our shipment from MIT is in the UK! currently
    • 7 Boxes will be de-containerised
    • Customs clearence is progressing as expected
    • Determining better path
      • UPS delivers all - We unload the smaller boxes, vessel remains outside and covered untill Kingsbury can move the vessel into T111
      • Kingsbury brings it all at once
      • Combination of the above
    • 27th is the preliminary show up date. Anyone with a gym membership can have a free workout if they want.
  • T111
    • Refurbishment is almost complete
    • Fume hood is operational - needs cleaning
    • Ethernet and Phone need connecting
    • Gas delivery into the room is a work in progress, still assessing the options
    • Compressed Air from next door - can come through the wall
    • Need to figure out how much storage we require along with work spaces.
    • Softwalled Cleantent to be installed early April.
    • GO GO GO


  • Long story short... Close, but no cigar

  • Things that have been done/tested
  • Drift field voltage is limited by the feedthrough, I get arcing at 50 Torr CF4 along the white lines
  • Carbon sputtered on the alumina section eventually causes a connection from the pin to ground
  • Rebuilt Everything
  • Plan is to apply Kapton directly to the inside hole and the inner metal cap with the goal of providing extra separation.
  • Current voltage records - Anode = 650V, Cathode 1000V
  • Translates Anode = 13,000 V/cm Drift field approx = 60 V/cm
  • Last weeks breakdown spots appear directly correlated to the source activity which was producing lots of local ionisation
  • Root_of_Evil_1.png

  • IMG_20170614_112634.jpg:
Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng GasCabinet.png r1 manage 179.8 K 15 Mar 2017 - 11:14 MarkWard  
PDFpdf HPTPC-UKHK.key.pdf r1 manage 14213.0 K 24 Apr 2017 - 13:27 MarkWard  
JPEGjpg IMG_0336.JPG r1 manage 786.0 K 29 Mar 2017 - 13:11 MarkWard  
JPEGjpg IMG_0342.JPG r1 manage 182.7 K 29 Mar 2017 - 13:11 MarkWard  
JPEGjpg IMG_20170327_094216.jpg r1 manage 2129.7 K 29 Mar 2017 - 13:10 MarkWard  
JPEGjpg IMG_20170526_170335.jpg r1 manage 1852.0 K 31 May 2017 - 12:22 MarkWard  
JPEGjpg IMG_20170614_112634.jpg r1 manage 1860.1 K 14 Jun 2017 - 13:03 MarkWard  
PNGpng Image-1.png r1 manage 1984.9 K 04 May 2017 - 10:24 MarkWard  
PNGpng O2Conc.png r1 manage 90.4 K 29 Mar 2017 - 12:33 MarkWard  
PNGpng O2Conc_trend.png r1 manage 149.1 K 29 Mar 2017 - 12:33 MarkWard  
PNGpng O2_Conc_HudAndFan.png r1 manage 44.9 K 05 Apr 2017 - 13:34 MarkWard  
PNGpng O2_Conc_HudOnly.png r1 manage 44.3 K 05 Apr 2017 - 13:33 MarkWard  
PNGpng Root_of_Evil_1.png r1 manage 40.6 K 08 Mar 2017 - 12:44 MarkWard  
PNGpng T111.png r1 manage 238.7 K 29 Mar 2017 - 12:52 MarkWard  
PNGpng T111_wHPTPC.png r1 manage 200.6 K 29 Mar 2017 - 13:04 MarkWard  
PNGpng Untitled.png r1 manage 204.3 K 10 May 2017 - 12:08 MarkWard  
PNGpng frame1.png r1 manage 130.8 K 10 May 2017 - 12:27 MarkWard  
PNGpng nd280.png r1 manage 288.3 K 22 Mar 2017 - 13:29 MarkWard  
Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r16 - 28 Jun 2017 - MarkWard

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback